OpenStudy (anonymous):

Can someone please show the work on how to solve this problem? Directions for the question: - Explain what the solution tells us about the lines represented by the equations. 3x - 4 y = 2 - x + 3y = 1

OpenStudy (anonymous):

Substitution or Linear Combination (Elimination) but id like to see the steps on how to do it .

OpenStudy (anonymous):

Start by multiplying the second equation by 3 and then adding the new equation to the first one.

OpenStudy (anonymous):

3x - 4y = 2 -3x +9y = 3 ------------ 0 + 5y = 5 y = 1

OpenStudy (anonymous):

3x - 4(1) = 2 3x = 6 x = 2

OpenStudy (anonymous):

thanks!